OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the given point and has the given slope. 1. (-6,-6) m=12 2. (-9,7) m=-1 3. (3,-11) m=0 4. (0.5,-1.5) m=2

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (♪chibiterasu):

Ah nvm, well the slope-int equation is y = mx + b

OpenStudy (♪chibiterasu):

@footballa That is point-slope form.

OpenStudy (anonymous):

But that's the formula you use to get it to slope intercept

OpenStudy (♪chibiterasu):

Mm, ok I see. Well let's continue.

OpenStudy (anonymous):

(-6,-6) m= 12. y- (-6) = 12( x-(-6)). y+6 = 12(x +6) y+6 = 12x + 72 y = 12x + 65

OpenStudy (anonymous):

OMG THANK YOU GUYS

OpenStudy (♪chibiterasu):

(-9,7) m=-1 (y - 7) = -1(x -9) (y - 7) = -1x + 9 y = -1x + 16

OpenStudy (♪chibiterasu):

(3,-11) m=0 (y - 3) = 0(x - 11) (y - 3) = 0 y = 3

OpenStudy (♪chibiterasu):

(0.5,-1.5) m=2 (y - 0.5) = 2(x + 1.5) (y - 0.5) = 2x + 3 y = 2x + 3.5

OpenStudy (♪chibiterasu):

That should be it.

OpenStudy (anonymous):

i love you.

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (♪chibiterasu):

Haha, no problem ;3